Solve by substitution method.
-10x+8y=-6
y=5/4x
I need help refreshing for final.

Answer by jim_thompson5910(35256)

if y=5/4x, then


-10x+8y=-6

-10x+8(5/4x)=-6


Notice how i replaced 'y' with '5/4x'. This is possible since 'y' and '5/4x' are the same


From here, solve -10x+8(5/4x)=-6 for x


Once you have a solution for x, use it to find y.